A chemist has 6 L of a solution that is 8% salt, How much water should be evaporated so that the solution that remains is 40% water?
Answer by Paul(988)

6(8)=40(6-x)
48=240-40x
-40x=-192
x=4.8

Hence, about 4.8L should evaporate.
Paul.